OpenStudy (anonymous):

1.1 metersx______=43.30 inches

OpenStudy (anonymous):

Oh I see. Didn't notice.

OpenStudy (anonymous):

1 meter = 39.37 inches

OpenStudy (mathstudent55):

\(1.1 ~ meters \times \dfrac{100~cm}{1~m} \times \dfrac{1 ~in.}{2.54~cm} = 43.30 ~inches \)